OpenStudy (anonymous):

Solve the equation by factoring: x2 - 7x + 12 = 0 A) x = 3 or 4 B) x = -3 or 4 C) x = 3 or -4 D) x = -3 or -4

OpenStudy (anonymous):

please don't just give me the answer.... I want to actually try to learn how to do it. Also, know how to do this the "plug and play" method, but I need to know how to do it when it's not multiple choice. Thanks!

OpenStudy (jiteshmeghwal9):

sorry dude

OpenStudy (anonymous):

why?

OpenStudy (jiteshmeghwal9):

first factorize the L.H.S. & then gt ur answer

OpenStudy (anonymous):

LHS?

Parth (parthkohli):

Hi, Sunshine! Would you like to know the whole process from the start?

OpenStudy (anonymous):

yes please! Thanks Spiderman! haha

OpenStudy (anonymous):

Sum = -7 Product = 12 two numbers -4 and -3 x^2 -4x - 3x + 12 =0 x(x-4) -3(x-4) (x-4) (x-3) x = 3 or x = 4

Parth (parthkohli):

All right then. Do you know what the standard form of a quadratic equation is?

OpenStudy (anonymous):

@Sunshine447 Did u understand!

OpenStudy (anonymous):

@ParthKohli yes

Parth (parthkohli):

Okay, what is it?

OpenStudy (anonymous):

ax + bx + c = 0 right?

Parth (parthkohli):

\(ax^2 + bx + c = 0\)* What you first do is find the product of \(a\) and \(c\). What are \(a\) and \(c\) here?

OpenStudy (anonymous):

@Sunshine447 Did u understand my method????

OpenStudy (anonymous):

a=1 and c=12

Parth (parthkohli):

Right! What's the product of 1, 12?

OpenStudy (anonymous):

12

Parth (parthkohli):

Correct! :) Now you have to find two numbers that add to get \(b\) and multiply to get the product of \(a\) and \(c\).

Parth (parthkohli):

Do you know two numbers such that: Their sum is -7. Their product is 12. ?

OpenStudy (anonymous):

no

Parth (parthkohli):

Hmm, you may find them by trial and error. Note that a negative number times a negative number is a positive number, so those numbers must be negative if the sum of negative.

OpenStudy (anonymous):

-3 and -4

Parth (parthkohli):

You got it! Now, we will 'split' the middle term with these coefficients. \(-7x = -3x - 4x \Longrightarrow x^2 - 7x + 12 = x^2 - 3x - 4x + 12\)

Parth (parthkohli):

Do you get it till this point?

OpenStudy (anonymous):

yeah i think so

Parth (parthkohli):

Okay, now we basically factor by 'grouping'. What you must do is factor the first two and last two terms. \(x^2 - 3x = x(x - 3)\) \(-4x + 12 = -4(x - 3)\) Get it till here?

OpenStudy (anonymous):

yeah

Parth (parthkohli):

So, \(x^2 - 3x - 4x + 12 = x(x - 3) - 4(x - 3)\). \(x(x - 3) - 4(x - 3) \Longrightarrow ( x - 4)( x - 3)\)

Parth (parthkohli):

Are you there? Do you get it?

OpenStudy (anonymous):

yeah

Parth (parthkohli):

Okay, so \((x - 4)(x - 3) = 0\). Do you know the 'zero-product rule'?

OpenStudy (anonymous):

i dont think so

Parth (parthkohli):

All right, you have two solutions here. \( \color{Black}{\Rightarrow x - 4 = 0}\) \( \color{Black}{\Rightarrow x - 3 = 0}\)

OpenStudy (anonymous):

okay...

Parth (parthkohli):

Can you solve those equations? That is how you get the solutions.

OpenStudy (anonymous):

x=4 and 3

OpenStudy (anonymous):

Thanks!

Parth (parthkohli):

You're welcome!

OpenStudy (theviper):

\[\LARGE{x^2-7x+12=0}\]\[\Large{x^2-3x-4x+12=0}\]\[\Large{x(x-3)-4(x-3)=0}\]\[\Large{(x-3)(x-4)}\]\[\Large{\color{green}{x=3\space or \space4}}\]

OpenStudy (theviper):

So A. is the right answer ;)

OpenStudy (theviper):

Understood ???